# FlexLLM: A System for Co-Serving Large Language Model Inference and Parameter-Efficient Finetuning

### v1

@@ 4 System Overview
Figure 4 shows an overview of FlexLLM, which co-serves PEFT and inference requests for LLMs. Users can choose to perform either finetuning or inference services on a model from the PEFT model hub, which stores the backbone LLM and all finetuned models. The programming interface for both inference and finetuning requests is unified through a PEFT-as-a-Service (PEFTaaS) interface. FlexLLM comprises three main system components. First, for finetuning requests that execute in parallel across multiple GPUs, the static compilation module (§5) generates a parallel computation graph. This graph specifies the execution of the PEFT model over the distributed environment and optimizes the graph by pruning unnecessary tensors for memory saving. Second, the dynamic scheduling module (§6) adapts a token-level finetuning mechanism. It mixes inference and finetuning tokens with a hybrid scheduling policy. Finally, the computation graph and schedule plan are executed by FlexLLM’s distributed co-serving runtime (§7).
Figure 5: Illustration of FlexLLM’s different dependent parallelization strategies with the example of LoRA.
@@ 4.1 PEFT-as-a-Service Interface
To facilitate the sharing of the backbone LLM between inference and finetuning, FlexLLM represents a PEFT model as a sequence of bypass networks attached to the backbone LLM. Each bypass network takes a single tensor from the backbone LLM as input and produces a single output tensor, which is added to one tensor of the backbone LLM. Let $X$ and $Y$ denote the input and output tensor of the bypass network, and let $f_{B}(X)$ and $f_{A}(X)$ denote the neural architecture of the backbone and bypass network for calculating $Y$. The bypass network can be formulated as $Y=f_{B}(X)+f_{A}(X)$. All existing PEFT methods can be represented in this format. For instance, the (IA)${}^{3}$ architecture modifies the topology of the backbone LLM by inserting an elementwise multiplication operator. It can be transformed into FlexLLM bypass network format since $Y=X\odot W=X+X\odot(W-O)$, where $\odot$ is elementwise multiplication and $O$ is a matrix of the same shape as $W$ whose elements are all one. A significant benefit of this interface is that all PEFT models preserve the neural architecture of the backbone LLM, which enables FlexLLM to fuse the computation graphs of different PEFT models. FlexLLM provides a unified user interface to launch inference and finetuning tasks by specifying a PEFT model and committing different types of input data, such as finetuning datasets or inference prompts.
@@ 5 Static Compilation
After a PEFT model is registered, FlexLLM compiles it into a parallel computation graph (PCG) [64]. To optimize the model’s finetuning across multiple devices, the system generates the best parallelization strategy by applying graph transformations to the PCG. Two key static compilation optimizations are dependent parallelization (Section 5.1) and graph pruning (Section 5.2). Dependent parallelization allows all PEFT models from the same backbone LLM architecture to share the backbone model parameters. The system generates the best parallelization plan for each PEFT based on the parallelization of the backbone LLM model. Graph pruning is another optimization technique that reduces the number of intermediate activation tensors that need to be saved during the forward pass and kept until the backward pass has completed. This technique helps to further reduce the memory requirements of finetuning the PEFT models. The rest of this section elaborates on the two techniques in depth.
@@ 5.1 Dependent Parallelization
To minimize memory overhead, FlexLLM enforces that all finetuning and inference tasks share the weights and intermediate activations of the base LLM, which requires the parallelization strategies of the PEFT models to be compatible with that of the base LLM, a task we called dependent parallelization.
To represent a parallelization strategy, FlexLLM generalizes the parallel computation graph representation introduced in Unity [64]. Similar to existing distributed LLM systems, FlexLLM supports parallelizing the base LLM using tensor-model parallelism, as introduced in Megatron-LM [59]. We first present Unity’s PCG representation and next describe how FlexLLM extends this representation. We also refer readers to [64, §3] for a more comprehensive description of PCGs.
Parallel computation graphs (PCGs).
PCGs leverage four parallelization primitives (i.e., partition, combine, replicate, and reduce) to collectively capture the communication pattern, data movement, and computation distribution associated with different parallelization strategies. Each tensor in a PCG is represented by a set of data dimensions, each of which includes two fields: a size and a degree. The degree field indicates the number of partitions the tensor is divided into along that dimension. Each tensor also includes a leading replica dimension, specifying the number of replicas of the tensor. A replica degree larger than 1 indicates that the tensor is replicated across multiple devices.
PCGs use four parallelization primitives to represent the computation and communication costs associated with different parallelization strategies. First, partition and combine change the parallelism degree of a tensor and are the “back propagation” of each other. In particular, partition increases the parallelism degree of a tensor dimension by equally splitting the dimension into multiple partitions, while combine reduces a tensor’s parallelism degree by concatenating multiple partitions into one. Second, replicate and reduce update the number of replicas of a tensor by copying and summing it. A more comprehensive introduction to the PCG primitives is available in the Unity paper [64, Figure 8].
Unity generates PCGs by iteratively applying graph substitutions on an input, unparalleled computation graph, which allows Unity to maintain the internal states of a tensor’s data dimensions implicit through the sequence of applied graphs substitutions. However, Unity’s approach requires starting from a fully unparalleled computation graph and does not work for dependent parallelization, where the base model has already been parallelized based on user-provided specification.
To address this issue, FlexLLM extends the PCG representation by introducing a state field for each dimension of a tensor. Each tensor dimension can be in one of the following four states, as shown in Figure 6. First, the non-parallel state indicates the tensor is not parallelized along this dimension (i.e., a parallelism degree of 1). Second, the partitioned and replicated states indicate the tensor is partitioned along the given dimension by splitting and replicating the original tensor. Finally, the pre-reduce state refers to the case where different replicas of the tensor each hold partial results, and a reduction across these replicas recovers the tensor’s final output. This scenario is common when parallelizing LLMs using tensor model parallelism (see an example for parallelizing the base model in Figure 5(a)).
Dependent parallelization problem.
Given multiple input and output parallel tensors and the parallelism states of their data dimensions, FlexLLM discovers a PCG that produces the output tensors with the given parallelism degrees and states by taking the input tensors as input.
Unlike the parallelization optimization problems addressed by existing auto-parallelization approaches, FlexLLM solves an optimization problem with additional constraints. Figure 5(a) shows the PCG for parallelizing the two linear layers using tensor model parallelism. Each green (or gray) box indicates a tensor algebra (or parallelization) operator, and each edge between operators represents a parallel tensor, whose dimensions are shown next to the parallel tensor. In this example, all parallel tensors have three dimensions, the first of which indicates the replica dimension and the remaining two correspond to the rows and columns of a matrix. We consider the dependent parallelization task for applying LoRA to the first linear operator, whose input and output tensors are parallelized in the [-, -, -] and [-,-,|] states, respectively. For the output tensor, the final partitioned state indicates the column of the output matrix is split across GPUs.
FlexLLM leverages the parallelism states of the input and output tensors for generating PCGs for the LoRA network. Specifically, FlexLLM inserts parallelization operators between tensor algebra operators and leverages the parallelism states of each tensor’s data dimensions to type-check the compatibility of these parallelization operators. Figure 6 shows how parallelization operators transit a tensor’s parallelism states. Most PEFT models only contain elementwise operators, which do not change the parallelism states of a tensor, and linear operators, which update an input tensor’s parallelism states by following the rules given by Figure 5(b).
To generate PCGs for a bypass network, FlexLLM inserts a parallelization operator between all tensor algebra operators, enumerates the type of the parallelization operator, and examines the compatibility of all tensors in the new PCG. A PCG that passes this examination is considered a candidate PCG for the dependent parallelization task. Figure 5(c) shows four candidates PCGs discovered by FlexLLM for parallelizing a LoRA network. These candidate PCGs introduce different parallelization operators and communication overheads. FlexLLM reuses the profiling-based cost model of Unity [64] and selects the candidate PCG with the lowest cost. For the four candidate PCGs in Figure 5(c), the first one achieves the lowest cost, since it does not introduce communications during forward or backward. Note that the replicate operator after the input tensor does not introduce cost since it can be fused with that of the base model. The other two parallelization operators introduce no communication cost since they are independent from the input and can be pre-processed.
@@ 5.2 Graph Pruning
This section introduces FlexLLM’s graph pruning algorithm, which takes a PEFT model and its backbone LLM as input and returns a minimal set of intermediate activations that must be reserved to finetune the given bypass networks. Graph pruning requires reasoning about the input and output tensors of each operator to prune nodes and edges, therefore FlexLLM introduces the following notations to the PCG representation. Let $\mathcal{G}=(\mathcal{N},\mathcal{E})$ denote the PCG of a PEFT model, where each node $n\in\mathcal{N}$ is a tensor algebra (or parallelization) operator and each edge $e=(n_{1},n_{2})\in\mathcal{E}$ is a tensor shared between operators. For each node $n$, let $\mathcal{I}(n)$ and $\mathcal{O}(n)$ denote its set of input and output tensors. Therefore, $(n_{1},n_{2})\in\mathcal{E}$ if and only if $\mathcal{O}(n_{1})\cap\mathcal{I}(n_{2})\neq\emptyset$ (i.e., $n_{2}$ depends on $n_{1}$).
@@
The design of FlexLLM’s graph pruning is based on two important observations. First, conventional ML training procedures maintain all intermediate activations in a forward pass for calculating gradients during the subsequent backward pass. However, due to the linear algebra nature of most operators, intermediate activations are reserved for computing the gradients for the trainable parameters of the base LLM, which are frozen during PEFT. This property allows FlexLLM to compute the gradients of the bypass networks of a PEFT model without using the majority of the intermediate activations. Second, different PEFT models attach bypass networks to different positions of the backbone LLMs and therefore require different intermediate activations for backpropagating through the bypass networks (see examples in Figure 8).
@@
Based on these observations, FlexLLM performs static graph pruning when generating the PCG for each registered PEFT model and dynamically schedules the finetuning and inference tasks of these PEFT models and allocates memory (see Section 6). Algorithm 1 shows FlexLLM’s graph pruning algorithm, which takes the PCG of a PEFT model $\mathcal{G}$ as input and generates an execution plan that specifies how to cache or recompute the intermediate activations during a forward pass for backpropagation. First, ReverseAutoDiff constructs the backward graph $\overline{\mathcal{G}}$ using the reverse-mode automatic differentiation algorithm [27]. Next, since all weights of the base LLM are frozen during finetuning, FlexLLM prunes gradients for these weights. Pruning an operator’s output tensors can enable subsequent pruning opportunities. FlexLLM maintains a queue $\mathcal{Q}$ to keep track of all operators whose input and/or output tensors have been updated and iteratively examines these operators to discover additional pruning opportunities. After graph pruning, FlexLLM generates a set of tensors $\mathcal{A}$ that contains all activations that need to be maintained for backward propagation. FlexLLM further combines graph pruning with two established memory optimizations: rematerialization and activation compression.
@@
In addition, FlexLLM also opportunistically applies lossless compression for remaining tensors whenever possible. This is because performing backward computation for some operators does not require access to the original input tensors. One example is the ReLU operator
$y=\texttt{ReLU}(x)=\begin{cases}x&x>0\\ 0&x\leq 0\end{cases}$
whose derivative is
$\frac{\partial y}{\partial x}=\begin{cases}1&x>0\\ 0&x\leq 0\end{cases}$
Therefore, instead of storing the original input tensor $x$, FlexLLM keeps the bitmask of $x$. Note that this compression does not introduce any accuracy loss.
@@ 6 Dynamic Scheduling
After static compilation generates a PEFT model’s executable generalized PCG, FlexLLM needs to determine the execution schedule to handle finetuning and inference requests concurrently. This is a challenge because of the distinct characteristics of each workload. For one, inference requests are received in a online fashion and are dynamically fluctuating, necessitating a system architecture that can rapidly adjust the GPU resources to ensure low latency. On the other hand, finetuning involves both a forward pass (like inference) and a backward pass to update model parameters. Third, the auto-regressive inference for generative LLMs requires a memory-intensive token-by-token generation pattern, while finetuning typically employs a compute-intensive sequence-level approach for better throughput. To address these challenges, FlexLLM employs a novel token-level finetuning mechanism and a hybrid token scheduler.
Figure 10: Illustration of KV gradients accumulation in the backward pass of FlexLLM’s token-level finetuning mechanism.
@@ 6.1 Token-Level Finetuning
To maximize GPU utilization, existing LLM finetuning systems generally process finetuning samples at the sequence level rather than the token level. As a result, directly batching inference requests with finetuning requests significantly increases the latency of each iteration and affects the SLO of inference requests.
We present our token-level finetuning mechanism in Algorithm 2. The main idea is to divide the finetuning process into smaller steps using a dynamic sliding window that consists of a series of tokens. This window is used for both the forward pass and the backward pass, and its size is determined by the hybrid token scheduler (line 4 and line 13), to ensure that the co-served inference tokens do not exceed the SLO constraints (§6.2). During the forward pass (lines 3-11), in each step, we slice the finetuning request into windows of $s_{i}$ tokens (line 5) and feed them into the model. The model processes the tokens layer by layer (lines 7-8), computes the generative loss of these tokens (line 10), and then moves to the next window of tokens (line 11). To maintain the same compute semantic as the sequence-level finetuning, we need to cache keys and values (similar to incremental decoding for inference [70]) as well as queries (reused by backward attention calculation) following the causal mask of autoregressive LLMs.
In the backward pass (lines 12-20), we get a slice of $s_{j}$ tokens’ loss in each step (line 14) and execute the model layers in reverse order to calculate the gradients (lines 15-16). However, unlike sequence-level attention execution, our token-level backward attention execution is more complex because of the auto-regressive token dependencies. We illustrate both the forward and backward PCGs of the attention module on the left and right sides of Figure 9. When we get a slice of $s_{j}$ tokens’ loss, the generated gradients of queries, keys, and values (i.e., $\Delta Q$, $\Delta K$, $\Delta V$) should have a shape of $[s_{j},h]$, where $h$ is the hidden size. The $\Delta Q$ calculation requires access to the key cache of all previous $l_{j}$ tokens, which aligns with the causal mask in forward execution. However, the autoregressive execution also makes the generated $\Delta V$ and $\Delta K$ be the size of $[l_{j},h]$ because the tokens’ loss within the window also contributes to all previous tokens. As shown in Figure 10, we propose to accumulate the KV gradients in each step and only apply the fully accumulated $s_{j}$ gradients (i.e., size of $[s_{j},h]$) to calculate the other model gradients until the backward traversal of all tokens in the sequence is completed.
@@
12:for $j\leftarrow 0,l_{0}\leftarrow L$; $l_{j}>0$; $j\texttt{++}$+ do $\triangleright$ Backward pass
13:     $s_{j}\leftarrow\textsc{HybridTokenScheduler}(l_{j},0)$
14:     $Y_{L-1,j}\leftarrow\textsc{Slice}(Loss,l_{j},s_{j})$
15:     for model layer index $n\in\textsc{Range}(N,0)$ do
16:         $Y_{n,j},\Delta K_{n,j},\Delta V_{n,j}\leftarrow\textsc{Backward}(M_{n},Y_{n-1,j},$
@@
18:         $\Delta KVAccum_{n}\leftarrow\textsc{Add}(\Delta K_{n,j},\Delta V_{n,j})$
19:         $G_{n,j}\leftarrow\textsc{Slice}(\Delta KVAccum_{n},l_{j},s_{j})$
20:         $\Delta M_{n}\leftarrow\Delta M_{n}+\textsc{CalculateGrads}(M_{n},G_{n,j})$
21:     $l_{j+1}\leftarrow l_{j}-s_{j}$
22:return $\Delta M$
Figure 11: Execution timeline and an example of token scheduling when co-serving inference requests and a finetuning mini-batch.

### v2

@@ 4 System Overview
Figure 2 shows an overview of FlexLLM, which co-serves both PEFT and inference requests for LLMs. Users can choose to use either finetuning or inference services on a model from the PEFT model hub, which stores the backbone LLM and all finetuned variants. The programming interface for both inference and finetuning requests is unified through a PEFT-as-a-Service (PaaS) interface. FlexLLM comprises three main components. First, for finetuning requests that execute in parallel across multiple GPUs, the static compilation module (§5) generates a parallel computation graph. This graph specifies the execution of the PEFT model over the distributed environment and optimizes the graph by pruning unnecessary tensors for memory saving. Second, the dynamic scheduling module (§6) adapts a token-level finetuning mechanism. It mixes inference and finetuning tokens with a hybrid scheduling policy. Finally, the computation graph and schedule plan are executed by FlexLLM’s distributed co-serving runtime (§7).
@@
Figure 4: Illustration of FlexLLM’s different dependent parallelization strategies with the LoRA example. Each green (or gray) box indicates a compute (or parallelization) operator, and each edge between operators represents a parallel tensor, and the parallelization states of the tensor’s dimensions are shown next to the edge.
@@ 4.1 PEFT-as-a-Service Interface
To facilitate the sharing of the backbone LLM across different tasks, FlexLLM represents a PEFT model as a sequence of bypass networks attached to the backbone LLM. Each bypass network takes a single tensor from the backbone LLM as input and produces a single output tensor, which is added to one tensor of the backbone LLM. Let $X$ and $Y$ denote the input and output tensor of the bypass network, and let $f_{B}(X)$ and $f_{A}(X)$ denote the neural architecture of the backbone and bypass network for calculating $Y$. The bypass network can be formulated as $Y=f_{B}(X)+f_{A}(X)$. All existing PEFT methods can be represented in this format. For instance, the (IA)3 architecture modifies the topology of the backbone LLM by inserting an elementwise multiplication operator. It can be transformed into FlexLLM’s bypass network format since $Y=X\odot W=X+X\odot(W-O)$, where $\odot$ is elementwise multiplication and $O$ is a matrix of the same shape as $W$ whose elements are all one. A significant benefit of this interface is that all PEFT models preserve the neural architecture of the backbone LLM, which enables FlexLLM to fuse the computation graphs of different PEFT models. FlexLLM provides a unified user interface to launch inference and finetuning tasks by specifying a PEFT model and committing different types of input data, such as finetuning datasets or inference prompts.
@@ 5 Static Compilation
Once a PEFT model is registered, FlexLLM compiles it into a parallel computation graph (PCG) [75]. To optimize finetuning performance, FlexLLM applies two key static compilation optimizations to discover an optimized PCG: dependent parallelization5.1) and graph pruning5.2).
@@ 5.1 Dependent Parallelization
To minimize memory overhead, FlexLLM allows finetuning and inference tasks to utilize the same backbone LLM and share memory allocated for both model weights and intermediate activations. This approach avoids redundant memory consumption while requiring the parallelization strategies of the PEFT models to be compatible with that of the backbone LLM, a task termed dependent parallelization.
Existing auto-parallelization approaches like Alpa [93] and Unity [75] can identify optimal parallelization strategies for backbone LLMs. FlexLLM extends the optimization target to also include the bypass networks in PEFT models. To achieve this goal, FlexLLM generalizes the parallel computation graph (PCG) abstraction introduced in Unity [75] by associating a state field with tensor dimension.
Figure 3 shows the four possible tensor states when parallelizing a tensor, including non-parallel (-), partitioned (|), replicated (=), and pre-reduce (+), as well as the state transitions. Given the fixed parallelization of the backbone LLM, FlexLLM searches for an optimal PCG for each bypass network by enumerating possible parallelization operators—partition, combine, replicate, and reduce—inferring output tensor states, and validating their compatibility. Figure 4 shows four candidate PCGs discovered by FlexLLM for parallelizing a LoRA network. To select the best strategy, FlexLLM reuses Unity’s profiling-based cost model [75] and chooses the candidate PCG with the lowest estimated execution cost.
@@ 5.2 Graph Pruning
FlexLLM’s graph pruning algorithm takes as input a PEFT model and its backbone LLM, and outputs a minimal set of intermediate activations that must be reserved to perform finetuning of the bypass networks. This pruning process requires reasoning over data dependencies between operators, specifically their input and output tensors. To facilitate this analysis, FlexLLM introduces additional notation in the PCG representation. Let $\mathcal{G}=(\mathcal{N},\mathcal{E})$ denote the PCG of a PEFT model, where each node $n\in\mathcal{N}$ represents a tensor algebra (or parallelization) operator, and each edge $e=(n_{1},n_{2})\in\mathcal{E}$ is a tensor shared between operators. For a node $n$, let $\mathcal{I}(n)$ and $\mathcal{O}(n)$ denote its set of input and output tensors. By definition, $(n_{1},n_{2})\in\mathcal{E}$ if and only if $\mathcal{O}(n_{1})\cap\mathcal{I}(n_{2})\neq\emptyset$, i.e., when at least one output tensor of $n_{1}$ serves as an input to $n_{2}$.
@@
The design of FlexLLM’s graph pruning algorithm is based on two key observations. First, conventional ML training procedures maintain all intermediate activations during the forward pass to compute gradients in the backward pass. However, due to the linear algebra nature of most operators, these intermediate activations are primarily used to compute gradients with respect to the base LLM’s trainable parameters. In the case of PEFT, the bacebone LLM parameters are frozen, and only the parameters in the bypass networks are updated. This insight allows FlexLLM to eliminate most intermediate activations, retaining only those necessary for computing gradients of the bypass networks. Second, different PEFT methods connect bypass networks at varying locations within the backbone LLMs, and therefore require different subsets of intermediate activations for backpropagation. Figure 6 illustrates examples of how various PEFT methods attach their bypass networks to the base LLM.
Building on these observations, FlexLLM performs static graph pruning during the construction of the PCG for each registered PEFT model, and dynamical scheduling to manage finetuning and inference tasks, along with corresponding memory allocations (see Section 6). Algorithm 1 shows FlexLLM’s graph pruning algorithm. Given a PEFT model’s PCG $\mathcal{G}$, the algorithm generates an execution plan that specifies whether each intermediate activation should be cached during the forward pass to enable efficient backpropagation.
@@
First, ReverseAutoDiff constructs the backward graph $\overline{\mathcal{G}}$ using reverse-mode automatic differentiation [33]. Since all weights of the base LLM are frozen during finetuning, FlexLLM prunes gradients the computation of their weights. Pruning the output tensors of an operator may, in turn, create opportunities to prune downstream operators. FlexLLM maintains a queue $\mathcal{Q}$ of operators whose input and/or output tensors have changed. It iteratively processes this queue to discover further pruning opportunities. Upon completion, FlexLLM identifies a set of tensors $\mathcal{A}$ representing the minimal set of intermediate activations that must be preserved for the backward pass. FlexLLM further combines graph pruning with two established memory optimizations: rematerialization and activation compression.
@@
In addition, FlexLLM also opportunistically applies lossless compression for remaining tensors whenever possible. This is because performing backward computation for some operators does not require access to the original input tensors. One example is the ReLU operator $y=\texttt{ReLU}(x)=\max{(x,0)}$ whose derivative is $\partial y/\partial x=1$ for $x>0$ and 00 for $x\leq 0$. Therefore, instead of storing the original input tensor $x$, FlexLLM keeps the bitmask of $x$.
@@ 6 Dynamic Scheduling
Once static compilation produces an executable generalized PCG for a PEFT model, FlexLLM must determine an execution schedule capable of concurrently handling both finetuning and inference requests. This is particularly challenging due to the distinct characteristics of these two workloads. First, inference requests arrive in an online and highly dynamic manner, requiring the system to rapidly adapt GPU resource allocation to maintain low-latency responses. In constrast, finetuning involves both a forward pass—similar to inference—and a backward pass to update model parameters. Additionally, auto-regressive inference for generative LLMs proceeds in a token-by-token fashion, while finetuning is typically performed at the sequence level to maximize throughput. To address these challenges, FlexLLM introduces a novel token-level finetuning mechanism along with a hybrid token scheduler.
Figure 8: Illustration of KV gradients accumulation in the backward pass of FlexLLM’s token-level finetuning mechanism, where $s_{j}$ and $l_{j}$ are the window length and the starting position of the $j$-th slice of the sequence.
@@ 6.1 Token-Level Finetuning
Processing a fine-tuning request typically involves performing both forward and backward passes across all model layers for the entire input sequence. When co-served with inference requests, this sequence-level execution can significantly interfere with the token-wise inference kernels and degrade system goodput—this is, the effective inference throughput that meets latency SLOs.
To address this problem and maximize GPU utilization, FlexLLM introduces a token-level finetuning mechanism, detailed in Algorithm 2. The key idea is to decompose the finetuning computation into smaller steps using a dynamic sliding window over tokens. This window spans both the forward and backward passes, with its size dynamically determined by the hybrid token scheduler (line 4 and 15) to ensure that inference requests continue to meet their SLO constraints (§6.2).
During the forward pass (lines 3-11), the finetuning sequence is partitioned into windows of $s_{i}$ tokens (line 5), which are incrementally fed into the model. The model processes each token window layer by layer (lines 7-8), computes the generative loss for the window (line 10), and then advances to the next set of tokens (line 11). To preserve the semantic of full sequence-level finetuning, FlexLLM caches key and value tensors—similar to incremental decoding in inference [84]—as well as query tensors, which are reused during backward attention computations. This caching follows the causal mask constraint inherent in LLMs.
During the backward pass (lines 12-21), FlexLLM executes the model layers in reverse order (line 13). Within each layer, the finetuning request is again divided into token windows (lines 14-16). However, unlike the relatively straightforward forward pass, token-level backward execution of the attention module is more complex due to auto-regressive token dependencies.
We illustrate the forward and backward PCGs of the attention module in Figure 7, with the forward graph on the left and the backward graph on the right. When computing gradients for a slice of $s_{j}$ tokens, the resulting gradients for the queries, keys, and values—denoted as $\Delta Q$, $\Delta K$, $\Delta V$)—have shapes of $[s_{j},h]$, where $h$ is the hidden size. To compute $\Delta Q$, the model must access the cached keys from all preceding $l_{j}$ tokens, consistent with the causal mask in forward pass.
However, due to the autoregressive nature of attention, the loss from the $s_{j}$ tokens also contributes to gradients with respect to previous tokens. As a result, both $\Delta K$ and $\Delta V$ have shapes of $[l_{j},h]$, encompassing all tokens prior to and including the current window. As shown in Figure 8, FlexLLM accumulates these KV gradients across steps. Only after the full backward traversal of the sequence is complete does FlexLLM apply the accumulated gradients (of size $[s_{j},h]$) to compute gradients for subsequent model parameters. Note that, this accumulation strategy only slightly increases memory consumption because FlexLLM’s layer-wise execution design allows each layer to reuse the workspace memory allocated by the previous layer, ensuring efficient memory utilization during backward pass.
@@
12:$Y_{N}\leftarrow Loss$
13:for model layer index $n\in\textsc{Range}(N-1,-1,-1)$ do
14:     for $j\leftarrow 0,l_{0}\leftarrow L$; $l_{j}>0$; $j\texttt{++}$+ do $\triangleright$ Backward pass
15:         $s_{j}\leftarrow\textsc{HybridTokenScheduler}(l_{j},0)$
16:         $Y_{n+1,j}\leftarrow\textsc{Slice}(Y_{n+1},l_{j},s_{j})$
17:         $Y_{n,j},\Delta K_{n,j},\Delta V_{n,j}\leftarrow\textsc{Backward}(M_{n},Y_{n+1,j},$
@@
19:         $\Delta KVAccum_{n}\leftarrow\textsc{Add}(\Delta K_{n,j},\Delta V_{n,j})$
20:         $G_{n,j}\leftarrow\textsc{Slice}(\Delta KVAccum_{n},l_{j},s_{j})$
21:         $\Delta M_{n}\leftarrow\Delta M_{n}+\textsc{CalculateGrads}(M_{n},G_{n,j})$
22:     $l_{j+1}\leftarrow l_{j}-s_{j}$
23:return $\Delta M$
Figure 9: Execution timeline and token scheduling when co-serving inference requests and a finetuning mini-batch.